OpenStudy (anonymous):

Does anyone know how to solve this for x? 4(x – b) = x

OpenStudy (diyadiya):

4x-4b=x 4x-x=4b 3x=4b x=(4/3)b

OpenStudy (anonymous):

Thank you so much!